OpenStudy (helpme1234):

If the volume of a prism is represented by the function x^3 + 11x^2 + 20x – 32. The width of the box is x – 1 while the height is x + 8. What is the expression representing length?????????? Please help and help me understand.

OpenStudy (helpme1234):

Anyone there???

OpenStudy (helpme1234):

Yay someone is typin

OpenStudy (anonymous):

v = whL x^3 + 11x^2 + 20x – 32 = (x-1)(x+8) x L the left side can be factored (x-1)(x+4)(x+8) = (x-1)(x+8)L then cancel (x+4) = L I think that is it.